# Comprehensive Review of Deep Unfolding Techniques for Next-Generation Wireless Communication Systems

### v2

@@ V-D IRS
To mitigate the blockage effect and extend the connectivity range for mmWave communication, IRS is a key technology proposed to enhance both energy and SE. IRS consists of many reflecting elements that steer the propagation of electromagnetic waves by adjusting their reflection coefficients. However, to unlock the full potential of IRS, accurate channel estimation is required for both individual and cascaded links. This is a challenging task because IRS-based channel estimation involves estimating multiple channels simultaneously: the direct channels between the BS and each user, the channel between the BS and IRS elements, and the channel between IRS elements and users.
In [109], Jiguang et al. proposed a method for estimating the rank-deficient cascaded channel in an IRS-assisted SIMO system. They developed a deep unfolding network based on a gradient descent algorithm. The model structure is inspired by DetNet [49]. In massive MIMO, increasing the number of IRS elements and BS antennas enhances the system’s capacity. However, this leads to increased training overhead to obtain accurate channel estimates within the channel coherence time. The proposed network addresses this challenge by compensating for the computational complexity.
Parallel factor decomposition is an efficient technique for estimating multiple large channel matrices in massive MIMO communication systems. It allows a high-dimensional matrix to be decomposed into a linear combination of several rank-one matrices, making the estimation of low-rank matrices more computationally efficient [182]. In [110], the parallel factor decomposition is unfolded to estimate multiple large channel matrices in the uplink of an IRS-empowered multi-user MISO communication system.
Awais et al. [111] developed two networks based on the MAP estimation criterion. The first network is a deep denoising network designed for noise-dominated SNR regions, providing support at various noise levels. The second network is a deep unfolded half-quadratic splitting network that alternates between optimizing extrapolation and denoising for interference-dominated regions. As the SNR increases, the proposed network shows significant improvement in normalized MSE. Under severe channel variations, the network delivers a performance gain of 20% in the high SNR range.
Tsai et al. [112] proposed a two-stage learned AMP network with row compression to address the joint estimation problem of direct and cascaded channels in IRS-aided mmWave systems. By utilizing the deep unfolding technique, the network combines the advantages of compressive sensing and deep learning. This approach enables reliable channel estimation with reduced training, which is crucial for efficiently reconfiguring wireless environments under IRS reflection control.
VI Precoder Design
This section reviews the applications of deep unfolding for precoder design used along with key 6G technologies.
VI-A Massive MIMO
In MIMO systems, beamforming enhances signal transmission and reception by directing signals toward specific directions using multiple antennas. This improves the received SNR [181, 183]. There are three types of beamformers: analog, digital, and hybrid. Fig.  6 illustrates the architectures for analog, digital, and hybrid beamforming [181]. The hybrid beamformer system model is described in [184]. A digital beamformer assigns one RF chain per antenna. However, in large-scale arrays, it has high power consumption and complex architecture. Analog beamformers use time delay units or phase shifters for antenna weighting but struggle with inter-user interference in multi-user scenarios. To overcome these issues, hybrid beamforming was introduced [185, 186].
Beamforming design is an optimization problem where the objective is to maximize a network utility function, such as sum rate, mean rate, or harmonic rate. This optimization is subject to resource and hardware constraints, including power budget and constant modulus analog beamforming coefficients.
Consider a base station (BS) with $n_{\text{t}}$ antennas serving $M$ single-antenna users [27]. The received signal at user $m$ is given by:
$y_{m}=\boldsymbol{{h}}_{m}^{H}\sum_{m=1}^{M}\boldsymbol{{f}}_{m}x_{m}+n_{m}$ (3)
where, $x_{m}$ is the transmitted signal with an average power of $\mathbb{E}\{|x_{m}|^{2}\}=1$, $\mathbf{h}_{m}\in\mathbb{C}^{n_{t}\times 1}$ is the channel vector, $n_{m}\sim\mathcal{CN}(0,\sigma^{2})$ represents the AWGN noise at user $m$, and $\boldsymbol{f}_{m}\in\mathbb{C}^{n_{t}\times 1}$ is the beamforming vector for user $m$. The signal-to-interference-plus-noise ratio (SINR) at user $m$ is given by:
$\text{SINR}_{m}=\frac{\left|{\boldsymbol{{h}}_{m}^{H}\boldsymbol{{f}}_{m}}\right|^{2}}{\sum_{i=1,i\neq m}^{M}\left|{\boldsymbol{{h}}_{i}^{H}\boldsymbol{{f}}_{i}}\right|^{2}+\sigma^{2}}.$ (4)
This expression quantifies the received signal quality by measuring the desired signal power relative to interference and noise. The hybrid beamformer design problem can be formulated as the maximization of the sum rate, given by:
$(\mathcal{P}_{\text{HBF}})\quad\max_{\boldsymbol{F}^{\text{DB}},\boldsymbol{F}^{\text{AB}}}$ $\sum_{m=1}^{M}\log_{2}\left(1+\text{SINR}_{m}\right)$ (5)
s.t. $\left\lVert\boldsymbol{F}^{\text{AB}}\boldsymbol{F}^{\text{DB}}\right\rVert_{F}^{2}\leq P,\quad\boldsymbol{F}^{\text{AB}}\in\mathcal{F},$
$|\boldsymbol{F}^{\text{AB}}|^{2}=1.$
where $\boldsymbol{f}_{m}=\mathbf{F}^{\text{AB}}\boldsymbol{f}_{m}^{\text{DB}}$, $\mathbf{F}^{\text{AB}}\in\mathbb{C}^{n_{t}\times N_{\text{RF}}}$ is the analog precoding matrix with $N_{\text{RF}}$ RF chains, belonging to a codebook $\mathcal{F}$. The digital precoding vector for the $m^{\text{th}}$ user is $\boldsymbol{f}_{m}^{\text{DB}}\in\mathbb{C}^{N_{\text{RF}}\times 1}$, and the digital precoding matrix $\boldsymbol{F}^{\text{DB}}$ consists of column vectors $\boldsymbol{f}_{m}^{\text{DB}}$. These problems are nonconvex and NP-hard, meaning they lack closed-form solutions. The unit-modulus constraints on the analog precoding matrix make precoder design even more challenging. To address this, the algorithm unrolling technique combines model-driven and data-driven approaches. It optimizes beamforming matrices by integrating the interpretability of model-based methods with the adaptability of data-driven techniques. Shlezinger et al. [26] discuss the architectural constraints and challenges of hybrid beamforming optimization using both conventional and data-driven methods.
The classic iterative WMMSE algorithm is widely used for precoder design to maximize the system’s weighted sum rate. It assigns different weights to the MSE of each user based on priority or quality of service. By adjusting these weights iteratively, WMMSE balances user performance and improves overall efficiency. The algorithm provides closed-form, iterative update rules to solve a surrogate optimization problem with the same local optima as the sum-rate maximization problem [122]. Maximizing the weighted sum rate while satisfying a power constraint is a nonconvex problem [187]. The approach proposed in [187] exploits the separable structure of power constraints. To achieve a locally optimal beamforming strategy, the authors introduce a simple and efficient distributed inexact cyclic coordinate descent method.
A deep unfolded version of the iterative WMMSE algorithm is proposed for precoder design in multi-user MIMO systems [69]. The authors call this network architecture the iterative algorithm induced deep unfolding network (IAIDNN). To improve efficiency, WMMSE is reformulated to approximate matrix inversions. A generalized chain rule in matrix form helps visualize the gradient recurrence between adjacent layers during backpropagation. Using this rule, the gradients of trainable parameters are computed effectively. The WMMSE algorithm [122] involves matrix inversions, eigendecompositions, and bisection searches in its iterations. Since any MMSE solution requires matrix inversion, WMMSE also relies on it. To improve convergence while avoiding matrix inversion, a downlink beamforming method using an unfolded PGD-WMMSE algorithm for multi-user MISO systems is proposed in [68]. The authors replace matrix inversions with PGD and compute gradients in parallel. Unlike IAIDNN [69], this approach eliminates matrix inversion, making hardware implementation more practical. Additionally, it achieves a high weighted sum rate in fully loaded scenarios, where the number of users matches the number of transmit antennas at the BS.
In [69], learnable parameters are used to approximate matrix multiplication and inversion in WMMSE update rules. However, the network fails to exploit the inherent graph structure of wireless networks. This limitation is addressed in [70]. An unfolded WMMSE algorithm for transceiver design in multicell MU-MIMO interference channels with local CSI is proposed in [70]. The authors use graph convolutional networks (GCNs) to unfold WMMSE, enabling coordination in multicell MU-MIMO systems. Graph filters and convolutional networks approximate certain WMMSE variables using graph learners. However, the approach in [70] lacks flexibility in learning the exact transformation of WMMSE variables. To overcome this, [71] introduces a hybrid unfolded WMMSE algorithm for transmit beamforming in MU-MIMO ad-hoc networks. This method leverages GNNs to learn functional transformations of WMMSE parameters. Compared to the classical WMMSE algorithm, the proposed network reduces the number of layers, leading to faster convergence of the optimization problem. In [188], the joint optimization of precoders for multicell MU-MIMO systems is studied. The WMMSE algorithm is unfolded to enable distributed and synchronized precoding across cooperative BSs. Numerical analysis compares this approach with the networks in [69, 68]. The methods in [69, 68] rely on centralized optimization, leading to high complexity and backhaul overhead. In contrast, the proposed network in [188] reduces computational cost while maintaining efficiency.
The gradient descent algorithm is a widely used method for finding a stationary point. Shi et al. [72] designed a hybrid analog-digital transceiver using iterative gradient descent to obtain minimum symbol error rate (MSER) in massive MIMO systems. They unfolded the iterative gradient descent algorithm into a DNN, where learnable parameters serve as step sizes for the gradients. By using a fixed number of DNN layers with optimized step sizes, the approach reduces the number of iterations and speeds up convergence compared to conventional methods. The input to this MSER-based algorithm includes a user’s transmit beamforming matrix, an analog transmit beamforming matrix, an analog receive beamforming matrix, and a digital receive beamforming matrix. The step sizes guide the gradient descent update for each matrix. The gradients for these parameters are updated using SGD. The authors claim that their proposed network matches the performance of the MSER-based iterative gradient descent algorithm while significantly reducing complexity. Zhu et al. [73] proposed a beamforming learning architecture by unfolding a parallel gradient projection algorithm for MISO heterogeneous systems. Their design optimizes scenarios where network parameters, such as the number of antennas, BSs, or overall network size, differ from those used during training. Gradient-based methods have been successfully applied to beamforming optimization with reduced complexity. However, these approaches can be further improved using advanced optimization frameworks like PGA.
A fast and robust precoding network for multi-user MIMO systems with imperfect CSI and per-antenna power constraints is proposed in [74]. The objective function is formulated as a geometric program that minimizes the MSE between the transmitted and recovered symbol vectors at the user. This optimization leverages the duality between downlink and uplink MSE feasible regions. The authors map the iterative optimization process into a layer-wise deep unfolding framework, enabling fast and robust downlink precoding prediction.
PGA is a variant of the gradient ascent algorithm used to maximize an objective function under constraints. It updates the solution in the gradient direction to increase the objective function and then projects it back onto the feasible set. Lavi et al. [75] unfolded the PGA algorithm to optimize hybrid precoding for multiuser downlink MIMO, assuming perfect CSI availability. To handle noisy CSI, they implemented the projected conceptual mirror prox algorithm [189, 190]. The step sizes for updating the analog and digital precoder matrices are tuned using mini-batch SGD, a data-driven approach. The solution is computed within a fixed number of iterations, defined as network layers, ensuring a balance between interpretability and efficiency. In [76], Bilbao et al. proposed an efficient analog beamforming design to maximize the achievable rate of massive MIMO in-band full-duplex systems. The beamforming problem is non-convex due to constant modulus constraints on analog beamforming coefficients. To address this, they introduced an unfolded algorithm combining PGA with ZF projection onto convex sets, achieving faster convergence.
Wang et al. [191] proposed a universal algorithm for downlink massive MU-MIMO systems with finite-alphabet precoding to minimize inter-user interference. The algorithm is an improved version of the original ADMM framework. The authors observed that the standard ADMM struggles with nonlinear discrete optimization due to biased alternating updates. To address this, they introduced iterative discrete estimation [191], ensuring unbiased updates and better convergence. To reduce hardware costs and power consumption in multi-user massive MIMO, He et al. [77] developed an unfolded network using the IDE-based algorithm with finite-alphabet precoding. Additionally, constructive interference-based symbol-level precoding for multi-user MISO downlink transmission is discussed in [78]. In [79], the conjugate gradient algorithm is unfolded for constant envelope precoding to reduce computational overhead. The Riemannian manifold optimization (MO)-based conjugate gradient descent is implemented across network layers. Simulations show that this approach effectively suppresses multi-user interference, improving overall performance.
In [192], the authors use deep unfolding to optimize user scheduling, BS clustering, and beamforming in a dense C-RAN system. Reducing the number of BSs per user helps ease backhaul congestion. The goal is to maximize the weighted sum rate while minimizing BS clustering, all within power constraints. The PGD algorithm is unfolded with step sizes learned in an unsupervised manner. This approach avoids matrix inversion and bisection search, reducing iterations and computational complexity.
In [80], the interior point method is unfolded in an unsupervised manner for precoding design. It leverages interference in multi-user MISO systems to minimize power while meeting SINR constraints. In [81], the authors address the challenges of mutual coupling in beamforming. They propose mitigating its effects through covariance matrix reconstruction by unfolding an augmented coprime array.
VI-B mmWave
mmWave wireless communication, operating in the 30–300 GHz band, offers abundant spectral resources to alleviate the bandwidth crunch at sub-6 GHz frequencies [157]. However, mmWave signals suffer from severe propagation and penetration losses. To compensate for these losses, large-scale antenna arrays with tens to hundreds of elements are required to achieve high power gains. The short wavelength (1–10 mm) of mmWave signals allows compact integration of these arrays into small transceivers. These features introduce unique challenges in designing physical-layer transmission algorithms [193]. mmWave MIMO systems rely on beamforming to mitigate multi-user interference while enhancing diversity and spectrum efficiency.
Typically, the beam selection and digital precoding matrices are designed separately, leading to suboptimal solutions. To address this, Hu et al. [86] proposed a joint design approach combining DL and model-driven techniques. DRL optimizes beam selection, while a deep unfolding neural network, based on iterative WMMSE, optimizes the digital precoding matrix. The goal is to maximize the system sum rate while satisfying beam selection constraints. The DRL and deep unfolding networks are trained sequentially, with one network’s output serving as the input to the other. Backpropagation and SGD are used for training and parameter updates.
Lin et al. [87] studied a linear MMSE transceiver design for uplink massive MU-MIMO with one-bit ADCs. To minimize the system MSE by optimizing the precoder at each user equipment, the PGD algorithm is unfolded. The authors focused on training only the step size in each PGD iteration, significantly reducing the training overhead. In [88], the authors combined deep generative models with unfolding techniques to achieve near-optimal hybrid beamforming while minimizing feedback and computational complexity. They represented the channel as a low-dimensional manifold using a GAN, enabling efficient optimization of digital and analog precoders. The hybrid beamforming problem is formulated as a bilevel optimization, where an iterative rule updates both precoders. The upper-level optimization refines the digital precoder, leveraging full knowledge of the lower-level task constraints. Simultaneously, the lower-level optimization adjusts the analog precoder based on the digital precoder’s decisions. This structured optimization approach ensures efficient beamforming design with reduced computational costs.
VI-C Next Generation Multiple Access
Rate-Splitting Multiple Access (RSMA) is a flexible multiple access technology that merges the strengths of space division multiple access (SDMA) and NOMA. It uses linearly precoded rate-splitting at the transmitter and SIC at the receivers. RSMA balances interference management by partially decoding some interference while treating the rest as noise, bridging the gap between SDMA (which treats all interference as noise) and NOMA (which fully decodes interference) [194]. In [107], the authors applied fractional programming with hyperplane fixed-point iteration to design the RSMA beamformer. The algorithm is unfolded into a neural network, replacing its inner loop with a small-scale DNN to predict Lagrangian dual variables. Simulations show that the proposed approach achieves high performance while reducing computational complexity.
VI-D IRS
An IRS consists of a programmable array of passive elements that dynamically adjust the phase shifts of incoming signals. This enables signals to align in phase at the receiver, enhancing reception. IRS can mitigate interference, create alternative propagation paths to bypass obstacles, and extend signal coverage to previously unreachable areas. The phase shifts are optimized to reflect signals coherently toward the intended user while minimizing interference in other directions. Therefore, jointly optimizing IRS phase control and transmit beamforming enhances the performance of IRS-assisted systems [195].
@@ VI-D IRS
With the aid of IRS, Chen et al. [113] addressed the joint design of the hybrid beamforming matrix and the phase-shifting matrix of the IRS elements. They proposed the deep unfolded WMMSE MO algorithm (DU-WMMSE-MO) using a deep unfolding methodology. The algorithm showed improved performance in IRS phase-shift design, utilizing two loops to jointly update the step size of the beamforming and IRS matrices. Compared to the Armijo backtracking line search method used in conventional WMMSE-MO [196], DU-WMMSE-MO provides more flexibility but with a higher runtime. Training is performed in an unsupervised manner, with a loss function that aligns with the objective function. The trainable parameters include the weights for the precoder, phase configuration of the IRS elements and combiner, and biases for the precoder and phase-shift matrix of the IRS.
@@
An IRS-assisted multiuser MIMO full-duplex system is explored in [114]. The active beamforming matrices at the BS and uplink users, along with the passive beamforming matrix at the IRS, are jointly optimized to maximize the weighted sum-rate of the system. To address the high computational complexity of matrix inversion in the conventional stochastic successive convex approximation (SSCA)-based optimization algorithm, a deep-unfolding neural network is proposed. Fig. 7 shows two such netwroks from [114] . The network consists of a long-term passive beamforming network (LPBN) and a short-term active beamforming network (SABN). The learnable parameters of the LPBN and SABN are the IRS phase shift or passive beamforming vector $\boldsymbol{\theta}$ and the active beamforming vector $\boldsymbol{\Psi}$, respectively.
Figure 7: Two networks LPBN and SABN [114].
@@
In [115], Ming-Min et al. introduced a two-timescale transmission approach to reduce the signal processing complexity and channel training overhead for the IRS-aided multiuser MISO system. To minimize the average transmit power at the AP while ensuring the individual quality of service constraints for the users, a primal-dual decomposition-based two-timescale hybrid beamforming method is proposed and unfolded. This approach allows for efficient beamforming optimization by separating the optimization into two timescales, improving both computational efficiency and performance.
@@
In [197], a low-complexity IRS-assisted downlink MISO system is designed using an unfolded block coordinate descent algorithm to address the joint optimization problem of transmit beamforming and IRS phase shifts. The multi-ratio fractional programming structure of the objective function is handled by decoupling the problem into two subproblems: one for optimizing the beamforming at the base station and another for optimizing the phase shifts of the IRS. This is achieved through the use of Lagrange dual transformation, effectively simplifying the optimization process. The numerical analysis validated the achieved weighted sum rate is comparable to [198] but with reduced computational complexity. Furthermore, the authors in [199], have developed an unfolded network to address the joint beamforming task for an IRS-assisted MISO system. The WMMSE and power iteration algorithm are used as an alternating optimization strategy to derive the local optimal solution for the active and passive beamforming. The proposed network reduce the computational runtime to approximately 3% of the block coordinate descent algorithm in [198].
In [116], an IRS-assisted MISO covert symbiotic radio communication system is explored, and a deep unfolding algorithm based on gradient descent is proposed for beamforming design. The goal is to optimize the achievable transmission rate of the primary transmitter by refining both active and passive beamforming strategies while satisfying two key constraints: (1) ensuring covertness and (2) maintaining the SNR of the secondary transmission. The complex optimization problem, along with its constraints, is transformed into the Lagrangian dual domain. Each layer in the deep unfolded network consists of four modules that update: (1) the active beamforming, (2) the passive beamforming, and (3) two Lagrange multipliers corresponding to the two constraints.
Gangyong et al. in [200] have designed the IRS phase shift matrix by maximizing the sum path gain of the channel. This maximization is achieved by unfolding the Riemannian gradient descent algorithm. The network takes channel matrices and the initial reflection coefficients as inputs and outputs the optimized IRS coefficients. By configuring the IRS element phase shifts, the system effectively directs signal propagation paths to maximize the sum-path-gain, which is a key metric for evaluating beamforming quality and learning optimal IRS phase shifts to enhance the overall beamforming pattern.
VI-E THz
@@
6G wireless networks are expected to achieve terabit-per-second (Tbps) data rates to support ultra-reliable low-latency communication. To enhance spectral and energy efficiency, THz-enabled massive MIMO has emerged as a promising solution, leveraging the abundant bandwidth at THz frequencies to significantly boost data rates. Additionally, hybrid beamforming offers a cost-effective approach by reducing the number of RF chains while maintaining high performance, making large-scale deployment of high-frequency systems more practical [201].
Nguyen et al. [118] proposed a deep unfolding framework by unrolling the iterations of Riemannian manifold optimization (Riemannian MO), specifically the MO-AltMin algorithm [202]. Instead of directly minimizing WMMSE and maximizing SE, the problem is transformed into an approximate matrix factorization problem, which is then solved using LS. This approach is computationally more efficient, as it requires fewer layers to compute the weights of the precoding matrix, significantly reducing the training time. Compared to the approach in [75], the proposed network achieves better performance while maintaining much lower time and computational complexity.
Minghui et al. [119] proposed a DL-based RSMA scheme for IRS-aided THz multi-user MIMO systems, focusing on enhancing robustness against CSI imperfections to improve system performance. The approximate WMMSE algorithm for RSMA is unfolded in this approach to boost precoding performance while reducing computational complexity. The deep unfolding algorithm processes the estimated equivalent channel through a series of layers to optimize the digital active precoding. This strategy aims to mitigate performance loss caused by model mismatches in real-world systems. By deep unfolding the approximate WMMSE algorithm, the method achieves improved precoding performance and lower computational complexity compared to the original WMMSE approach. This iterative deep unfolding process refines the precoding strategy based on the estimated equivalent channel, optimizing performance in IRS-aided THz massive MIMO-OFDM systems.
In [120], Adam et al. introduced a method to jointly optimize active beamforming, passive beamforming, and UAV trajectory to minimize power consumption while accounting for UAV jitters and imperfect hardware constraints. They designed a hybrid semi-unfolding DNN to address these constraints during the real-time CSI gain estimation for IRS and UAV links. The approach is termed “semi-unfolded” because it combines two types of networks: one based on deep unfolding and the other based on a multi-kernel convolutional LSTM (MK-CLSTM) NN. For active beamforming, the authors employed semidefinite programming to obtain a closed-form solution, which was then transformed into NN layers to replicate the iterative beamforming solution in real time. The Lagrangian functions with well-defined Karush-Kuhn-Tucker (KKT) conditions were derived to design the layers of the deep unfolded architecture. The UAV trajectory and beamforming vectors are the parameters updated within these layers. For passive beamforming, a MK-CLSTM-based NN was used. Training was done using mini-batch SGD. The authors claim that their model-driven approach achieved $99.24\%$ accuracy and is $26.21\%$ faster than the classical iterative methods, offering significant improvements in both efficiency and speed for UAV-based communication systems.
VII Sensing and Communication
The independent operation of sensing and communication as separate systems has been widely studied [203]. However, with the deployment of wireless communication services in higher frequency bands, such as mmWave and THz, spectrum congestion has become a major issue. In the 6G era, independent sensing systems experience interference from communication signals, leading researchers to recognize the shared system architecture and signal processing techniques between sensing and communication [204], [205], [206]. This realization has driven the development of ISAC, where both functionalities are merged into a unified system.
ISAC can be categorized based on sensing methods into active sensing and passive sensing [207].
  • Active sensing (Mono-static sensing): The transmitter and receiver are co-located, allowing the system to extract target parameters using both the transmitted signal and received echoes. However, this approach suffers from significant self-interference at the ISAC receiver due to the strong reflections from the transmitted signals.
  • Passive sensing (Bi-static sensing): The sensing echoes are received at a different location than the transmitter, eliminating the need for in-band full-duplex operation at the receiver. This approach is more suitable for applications where reducing self-interference is a priority.
ISAC provides a promising solution for optimizing spectral efficiency and enhancing system performance in future wireless networks. However, understanding the fundamental performance limits of ISAC systems is crucial for their practical implementation. The estimation accuracy of key sensing parameters, such as direction of arrival, Doppler frequency, velocity, and target position, was thoroughly investigated in [208]. This study provided a detailed analysis of both theoretical and practical limits of ISAC systems. The authors emphasized that the performance limits of ISAC channels cannot be derived by simply merging the existing bounds from independent sensing and communication systems. Instead, a unified framework is necessary to characterize these limits effectively. To address this, the authors proposed deriving unified Cramér-Rao bounds for ISAC, which establish the fundamental lower bounds on sensing accuracy given a specific allocation of orthogonal resources between communication and sensing. These unified expressions provide insights into how ISAC systems can optimize their resource allocation strategies to achieve superior performance in joint sensing and communication tasks.
VII-A MIMO ISAC
The integration of deep unfolding in ISAC systems has shown promising advancements in optimizing waveform design, beamforming, and communication-sensing tradeoffs. Ziwei et al. [82] applied deep unfolding techniques to optimize waveforms in MIMO radar systems, specifically focusing on minimizing the range-integrated sidelobe level. This is crucial for reducing interference and improving target detection performance. Their deep unfolding network efficiently optimizes radar waveforms, enhancing overall radar performance. In [83], the challenge of designing constant-modulus waveforms for multiuser MIMO ISAC systems was explored. Since constant-modulus waveforms help avoid signal distortion in nonlinear power amplifiers, their design is crucial. However, the non-convex nature of the constant-modulus constraint makes optimization challenging. To address this, the authors employed a deep unfolded network with a sparsely connected structure, trained in an unsupervised manner, ensuring efficient waveform design under these constraints. Hybrid beamforming in ISAC systems poses additional complexity due to the need to balance communication rates and sensing accuracy. The authors in [84] tackled this challenge by unfolding the iterative PGA algorithm. Their deep unfolding-based method accelerated convergence and enhanced overall system performance, achieving an optimal balance between communication and sensing objectives. Further, joint optimization of radar-communication beamforming was addressed in [209]. The authors incorporated symbol-level precoding with constructive interference constraints to improve energy efficiency by exploiting multi-user interference. They developed iterative algorithms to leverage problem separability, enabling parallel and recursive computation. Deep unfolding was utilized to reduce computational complexity, improve scalability, and support real-time implementation.
VII-B mmWave MIMO ISAC

### v3

@@ V-D IRS
With the aid of IRS, Chen et al. [171] addressed the joint design of the HBF matrix and the phase-shifting matrix of the IRS elements. The authors proposed the deep unfolded WMMSE-MO network (DU-WMMSE-MO), for IRS-assisted mmWave MIMO-OFDM systems. The trainable parameters include the weights for the precoder, phase configuration of the IRS elements and the combiner, and biases for the precoder and phase-shift matrix of the IRS. The algorithm showed improved performance in IRS phase-shift design, utilizing two loops to jointly update the step-size of the beamforming and IRS matrices. Compared to the Armijo backtracking line search method used in classical WMMSE-MO [172], DU-WMMSE-MO [171] provides more flexibility but with a higher runtime.
@@
An IRS-assisted multiuser MIMO full-duplex system is explored in [173]. The active beamforming matrices at the BS and uplink users, along with the passive beamforming matrix at the IRS, are jointly optimized to maximize the weighted sum rate of the system. A deep unfolding approach is proposed to approximate the stochastic SSCA-based beamforming algorithm, reducing the complexity of matrix inversion while maintaining near-optimal performance.
@@
In [174], Min et al. introduced a two-timescale transmission approach to reduce the signal processing complexity and channel training overhead for the IRS-aided multiuser MISO system. To minimize the average transmit power at the analog precoder while ensuring the individual quality of service (QoS) constraints for the users, a primal-dual decomposition-based two-timescale HBF method is proposed and unfolded. This approach allows for efficient beamforming optimization by separating the optimization into two timescales, improving both computational efficiency and performance.
@@
In [175], a low-complexity IRS-assisted downlink MISO system is designed using an unfolded block CD algorithm to address the joint optimization problem of transmit beamforming and IRS phase shifts. The multi-ratio fractional programming structure of the objective function is handled by decoupling the problem into two subproblems: one for optimizing the beamforming at the BS and another for optimizing the phase shifts of the IRS. This is achieved through the use of Lagrange dual transformation, effectively simplifying the optimization process. The numerical analysis validated that the achieved weighted sum rate is comparable to that of [176] but with reduced computational complexity. Furthermore, the authors in [177], have developed an unfolded network to address the joint beamforming task for an IRS-assisted MISO system. The WMMSE and power iteration algorithm are used as an alternating optimization (AO) strategy to derive the local optimal solution for the active and passive beamforming. The proposed network reduce the computational runtime to approximately 3% of the block CD algorithm in [176].
In [178], an IRS-assisted MISO covert symbiotic radio communication system is explored, and a deep unfolding algorithm based on gradient descent is proposed for beamforming design. The goal is to jointly optimize active and passive beamforming under covert communication constraints. The step-size for updating the active and passive beamformers, and Lagrange multipliers form the trainable set. In contrast, Gangyong et al. in [179] focused purely on passive beamforming and unfolded the Riemannian gradient descent algorithm. They have designed the IRS phase shift matrix by maximizing the sum path gain of the channel. In this network, the only trainable parameter is the step-size associated with each Riemannian gradient descent iteration. The network in [179] is simpler compared to that of [178], due to less number of parameters to train but is specific to optimizing the phase shift of the IRS.
V-E Terahertz
@@
6G wireless networks are expected to achieve terabit-per-second (Tbps) data rates to support ultra-reliable low-latency communication. To enhance spectral and energy efficiency, terahertz (THz)-enabled massive MIMO is a promising solution, leveraging the abundant bandwidth at THz frequencies to significantly boost data rates [180].
Nguyen et al. [181] proposed a deep unfolding framework that unfolds the iterations of Riemannian MO (specifically the MO-AltMin algorithm) [182]. Instead of directly solving the original SE maximization or WMMSE minimization problems, which are highly nonconvex and computationally intensive, they reformulated the HBF design as an approximate matrix factorization problem. This factorization leads to LS formulation for analog and digital beamforming enabling a more efficient solution. By leveraging this structure, their proposed ManNet-based approach [181] requires significantly fewer layers to estimate the analog precoder, resulting in faster training and lower computational complexity as compared to the approach in [155].
Minghui et al. [183] proposed an unfolded network to enhance RSMA precoding in IRS-aided THz MU-MIMO systems under CSI imperfections. They unfold an approximate WMMSE algorithm to optimize the digital active precoding layers based on the estimated equivalent channels. The trainable set includes the transformer weights associated with the private and common streams and a regularization parameter. Compared to the original approximate WMMSE algorithm, the proposed unfolding method achieves superior precoding performance with substantially lower computational complexity.
V-F Discussions and Takeaways
Precoding design in massive MIMO and HBF systems involves solving non-convex optimization problems under hardware constraints and channel uncertainty. Popular unfolded networks for precoding design are listed in Table V. For instance, IAIDNN [146] unfolds the WMMSE algorithm and replaces complex operations like matrix inversion with a learnable block. This results in a significant reduction in computational complexity while achieving performance close to traditional WMMSE solutions. Deep unfolding enables the incorporation of domain-specific constraints such as the constant modulus constraint in analog beamformers through tailored activation functions. This approach significantly reduces computational complexity by avoiding matrix inversion and accelerates convergence, making it well-suited for real-time hybrid precoding under practical hardware limitations.
Despite their benefits, current unfolded precoding models often assume fixed system dimensions and ideal CSI, which limits generalization and applications in practical systems. Future work should focus on scalable, adaptive unfolding frameworks that can operate reliably in time-varying, multi-user, and wideband settings. Integration with channel estimation modules and robust designs under CSI uncertainty will be key for practical deployment in next-generation wireless networks. Moreover, applying such unfolded frameworks for precoders in real systems and analyzing the performance in real experiments will be very beneficial.
VI Sensing and Communication
The independent operation of sensing and communication faces challenges in the 6G era due to spectrum congestion in high-frequency bands like mmWave and THz [184]. This has driven the development of ISAC systems, which unify both functions using shared architectures and processing techniques [185], [186], [187]. ISAC includes active (mono-static) and passive (bi-static) sensing methods [188]. While ISAC enhances SE and system performance, understanding its fundamental limits is crucial. Key sensing parameter estimations have been thoroughly analyzed in [189], revealing theoretical and practical boundaries.
TABLE VI: Summary of typical deep unfolding models for ISAC and their advantages and disadvantages.
Unfolding Model Main Novelties and Performance Gains Advantages and Disadvantages
Unfolded PGA network [191] Unfolds PGA to maximize the system sum rate while minimizing the sensing beampattern error. Achieves $33.5\%$ higher sum rate, $2.5$ dB lower radar beam pattern error compared with SCA, and $65\%$ reduction in runtime and complexity compared with the conventional PGA. Advantages: Ensure convergence in the challenging ISAC HBF designs, perform well in both single use and MU systems. Disadvantages: The convergence speed is relatively slow, high complexity to compute the gradients.
ISAC-Net [194] Unfolds the denoising AMP method for communication and DFT-based sensing in narrowband OFDM ISAC. Achieve communication performance close to OAMP-Net2 [11] and better sensing performance than classical 2D-DFT algorithm. Advantages: Leverages shared resources to overcome challenges in signal processing without prior knowledge of transmitted signals, enhanced accuracy and security. Disadvantages: Beam squint effect is not considered leading to discrepancies extracting sensing parameters (angle of arrival or angle of departure) across various subcarriers in OFDM.
ADMM-Net for OTFS ISAC [196] Unfolds ADMM to estimate delay-Doppler targets and demodulate OTFS symbols under high mobility. With $2$ targets, ADMM-Net achieve $4$ dB gain in NMSE compared with the classical ADMM and ISTA at SNR=$14$ dB. Advantages: Suitable for vehicular/high-mobility scenarios and robust to delay-Doppler spread. Disadvantages: Real time CSI estimation is not addressed.
Double loop unfolded network [197] Unfolds gradient descent to jointly optimize IRS phase shifts in ISAC system. Minimizes worst case CRB for sensing targets. Advantages: Has low computational complexity and fast convergence (within $15$ layers). Disadvantages: Offline training ($1000$ channel realization, $20$ epochs) may not be feasible for frequently changing environments.
VI-A MIMO ISAC
The integration of deep unfolding in ISAC systems has shown promising advances in optimizing waveform design, beamforming, and communication-sensing trade-offs. In [190], the challenge of designing constant-modulus waveforms for MU-MIMO ISAC systems was explored. Since constant-modulus waveforms help avoid signal distortion in non-linear power amplifiers, their design is crucial. However, the non-convex nature of the constant-modulus constraint makes optimization challenging. To address this, the authors employed a deep unfolded network based on PGD algorithm, ensuring efficient waveform design under these constraints.
HBF in MIMO ISAC systems poses additional complexity due to balancing communication rates and sensing accuracy. The authors in [191] tackled this challenge by unfolding the iterative PGA algorithm. Their deep unfolding-based method accelerated convergence and enhanced overall system performance, achieving an optimal balance between communication and sensing objectives. In addition, joint optimization of radar communication beamforming in MIMO system was addressed in [192]. The authors incorporated symbol-level precoding with constructive interference constraints to improve energy efficiency by exploiting multi-user interference. Since, the objective problem is computationally complex due to non-convexity and higher dimensional matrix inversions, a custom iterative algorithm is developed. The iterative algorithm is based on block CD and primal-dual optimization (augmented Lagrangian method). Therefore, the trainable parameter set for training the deep unfolded network include step-sizes to update the primal, dual and auxiliary variable, as well as a correction matrix to approximate matrix inversion. The proposed network reduces computational complexity due to small number of trainable parameters, improve scalability, and support real-time implementation.